# Flipping the Dialogue: Training and Evaluating User Language Models

### v1

@@ 6 Related Work
Historically, one of the main uses of computers has been to simulate environments across many domains, from video games, macro-economics Holland & Miller (1991); Zheng et al. (2022), weather and climate Taylor et al. (2012), political elections Straffin Jr (1980); Le Pennec & Pons (2023), biological simulators for molecular interaction Shaw et al. (2010), financial markets Cont & Bouchaud (2000), self-driving car environments Dosovitskiy et al. (2017), and physics simulators for robotics Todorov et al. (2012). Simulation environments are invaluable in these applications as they enable the evaluation of theories or methods, on-the-fly decision-making for a system based on simulation, or the training of AI systems, for instance through reinforcement learning. As conversational interfaces see widespread adoption in society, we view UserLMs as a foundational simulation environment for user-AI conversation.
@@
There is relatively much less work on how to evaluate simulators at replicating human behavior in their conversations with assistants Deriu et al. (2021). Some studies evaluate simulators by measuring the correlation between their utterances and human utterances using lexical or stylistic metrics Ivey et al. (2024); Pietquin & Hastie (2013), or the diversity of their generated utterances Wang et al. (2025). Others have resorted to human evaluation, asking crowd-workers to judge the quality of the simulated conversations Kranti et al. (2025); Sekulić et al. (2024; 2022); Shi et al. (2019). Another line of work performs a more indirect evaluation of the simulator by examining if the task-solving performance when interacting with the assistant correlates with the performance from interactions with real users Chang et al. (2025). Our study introduces fine-grained evaluations of key simulator properties, focusing on multi-turn interaction behavior and robustness. We also assess simulators when deployed for solving coding and math tasks with assistants, analyzing aspects such as information coverage, conversational pace, and diversity.
@@ Acknowledgments
We would like to thank Chan Young Park, Jillian Fisher, and Serina Chang, and the AI Interaction and Learning team members at Microsoft Research for valuable discussions. We also thank Pangram for providing credits to access their API, facilitating some of our robustness evaluations.
@@ Appendix A Additional Experimental Details
During initial analysis, we noticed a lot of near-duplicate conversations in WildChat - often consisting of prompt templates with minor input variations. To improve the quality of the training corpus, we applied a 7-gram counter on the first-turn user prompts and manually examined the most frequent patterns. We identified several recurring prompt templates, with one used in as many as 81,196 conversations, that do not represent natural user interaction and could degrade model quality. Removing all near-duplicates reduced the dataset to 384,336 conversations.

### v2

@@ 6 Related Work
Historically, one of the main uses of computers has been to simulate environments across many domains from macro-economics Holland & Miller (1991); Zheng et al. (2022), climate Taylor et al. (2012), elections Straffin Jr (1980); Le Pennec & Pons (2023), molecular biology Shaw et al. (2010), financial markets Cont & Bouchaud (2000), self-driving cars Dosovitskiy et al. (2017), and robotics Todorov et al. (2012). Simulation environments are invaluable in these applications as they enable the evaluation of theories or methods, on-the-fly decision-making for a system based on simulation, or the training of AI systems, for instance through reinforcement learning. As conversational interfaces see widespread adoption in society, we view user LMs as a foundational simulation environment for user-AI conversation.
@@
There is little work on how to evaluate simulators at replicating human behavior in their conversations with assistants Deriu et al. (2021). Some studies measure the correlation between simulator and human utterances using lexical or stylistic metrics Ivey et al. (2024); Pietquin & Hastie (2013), or the diversity of their generated utterances Wang et al. (2025). Others have resorted to human annotation to judge the quality of the simulated conversations Kranti et al. (2025); Sekulić et al. (2024; 2022); Shi et al. (2019). Another line of work evaluates the simulator more indirectly by testing whether task performance with the assistant correlates with performance in real user interactions Chang et al. (2025). Our study introduces fine-grained evaluations of simulator properties related to multi-turn interaction behavior and robustness. We also assess simulators when deployed for solving coding and math tasks with assistants.
@@ Acknowledgments
We would like to thank Chan Young Park, Jillian Fisher, and Serina Chang, and the AI Interaction and Learning team members at Microsoft Research for valuable discussions. We also thank Pangram for providing credits to access their API, facilitating some of our robustness evaluations. Tarek Naous and Wei Xu are partially supported by the NSF CAREER Award IIS-2144493. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Additional Experimental Details
During initial analysis, we noticed a lot of near-duplicate conversations in WildChat - often consisting of prompt templates with minor input variations. To improve the quality of the training corpus, we applied a 7-gram counter on the first-turn user prompts and manually examined the most frequent patterns. We identified several recurring prompt templates, with one used in as many as 81,196 conversations, that do not represent natural user interaction and could degrade model quality. Removing all near-duplicates reduced the dataset to 384,336 conversations. To verify the efficacy of our filtering and data splitting, we computed the 7-gram overlap between the first-turn user prompts in the train and test sets, finding only 4.19% overlap. This low overlap confirms that our deduplication and user-based splitting procedure effectively minimizes data leakage between splits.